# Limits to scalable evaluation at the frontier: LLM as Judge won’t beat twice the data

### v1

@@ 1.1 Our contributions
Figure 1: Model ranks on MMLU in a semi-synthetic setting (a) and for the top-10 models on HELM by July 2024 (b). Ranks are heavily perturbed despite high judge accuracy.
@@
  • Our main result (Theorem 6) then shows that this correlation is small whenever the judge model performs worse than the evaluated model. In that case the best possible outcome is an effective doubling of the number of samples $s$, no matter how many samples of the proxy score $\tilde{s}$ are available (Corollary 7).
@@
In addition, we show that the commonly reported agreement rate between the judge and the ground truth neither meaningfully restricts the bias introduced by the judge model, nor the sample size savings for debiasing methods. Instead, we propose a novel balanced agreement metric that provides meaningful bounds for sample size savings (Theorem 9).
@@ 2 Formal setup and motivation
In this section, we introduce the formal model of evaluation that we work with throughout. We then show how judge bias threatens the validity of LLM-based evaluations, even if judge models are quite accurate.
@@
2.1 Judge bias can strongly perturb rankings
The goal is now to understand how to employ the proxy score $\tilde{s}$ to benchmark a set of models $\mathcal{M}$ in a valid way. Following the reasoning in Salaudeen & Hardt (2024), benchmarks have two purposes: quantifying model performance and ranking models. In turn, a useful LLM judge should accurately estimate each model’s performance, as well as induce the same ranking of evaluated models as the ground truth. For the first goal, the main quantity of interest is the judge bias , defined as
@@ 2.2 High agreement rate is insufficient for evaluation
We note that all judge models from the last section perform well at their task: LLama-3 70B has $79\%$ accuracy on MMLU, while GPT-4 has $84\%$, and the models used in MT-Bench have $85\%$ agreement with expert annotators on average. Our next proposition explains how it is possible for model rankings to be strongly perturbed despite accurate judges. Before stating it, we need to formally define the agreement $\operatorname*{\mathrm{AG}}(m)$ between score $s$ and proxy score $\tilde{s}$ as
@@ 3 Debiasing LLM Judgments
In the previous section, we demonstrated how biased LLM judges can lead to strongly perturbed model rankings. This motivates the need for debiasing methods. We begin by discussing a simple sufficient condition to guarantee asymptotically correct rankings: Ensure that our estimates of $\operatorname*{\mathbb{E}}s(m)$ are unbiased for all models $m$. Statistical bias correction methods can make use of a large number of proxy scores and a smaller number of gold standard labels to construct an unbiased estimator. We discuss one such method in detail, and show that it is essentially optimal in terms of estimator variance. Afterwards we prove our main result, i.e., that LLM judges offer limited benefits for evaluating state-of-the-art models. We conclude the section with additional upper and lower bounds on the benefits of judge labels based on a novel balanced agreement metric.
@@ 3.1 Uniformly controlling judge bias ensures correct rankings
In particular, this implies that if $|\operatorname*{\mathrm{JB}}(m)|\leq\frac{\epsilon}{2}$ for all models $m\in\mathcal{M}$, a sufficiently large amount of proxy samples $\tilde{s}$ will give us the correct rankings with high probability. However, it is important for the condition to hold for all models. Judge bias might be low averaged over models, even if the judge overestimates the performance of some models and underestimates the performance of others.
@@ 3.2 Background on Prediction Powered Inference
While Proposition 3 suggests an avenue to obtain correct model rankings when using a proxy score, actually ensuring a small judge bias $|\operatorname*{\mathrm{JB}}(m)|$ for all evaluated models would require near-perfect judge models. Instead, we switch our attention to recently proposed statistical debiasing methods that make use of a small amount of ground truth data to debias a given proxy score.
@@
As a linear combination of unbiased estimators this remains unbiased, but we can now optimize over $\lambda$ to minimize estimator variance. We obtain the optimum value
$\lambda^{*}=\frac{\operatorname*{\mathrm{Cov}}(s(m),\tilde{s}(m))}{\left(1+\frac{n}{N}\right)\operatorname*{\mathrm{Var}}\tilde{s}(m)}.$
The estimator $\hat{\theta}^{PP}_{\lambda^{*}}$never increases variance compared to the classical ground truth estimator:
@@
As the variance of $\hat{\theta}^{GT}$ scales as $\Theta(1/n)$, using an estimator $\hat{\theta}$ with $r=\frac{\operatorname*{\mathrm{Var}}\hat{\theta}}{\operatorname*{\mathrm{Var}}\hat{\theta}^{GT}}$ is equivalent to increasing the ground-truth sample size by a factor of $\tau(\hat{\theta})\coloneqq\frac{1}{r}.$ We call $\tau(\hat{\theta})$ the sample efficiency factor of the estimator $\hat{\theta}$. For fixed values of $n$ and $N$, the relative variance reduction of $\hat{\theta}^{PP}_{\lambda^{*}}$ compared to $\hat{\theta}^{GT}$ is determined by the squared Pearson correlation coefficient $\rho$ between the ground truth labels $s$ and model judgements $\tilde{s}$:
Therefore, we have
$\tau(\hat{\theta}^{PP}_{\lambda^{*}})=\frac{1}{1-\frac{1}{1+\frac{n}{N}}\rho(s(m),\tilde{s}(m))^{2}}\geq 1$
The sample efficiency factor $\tau(\hat{\theta}^{PP}_{\lambda^{*}})$ is large if the squared correlation between true and proxy scores is high. Conversely, if the squared correlation is small, so is the sample efficiency factor.
3.3 PPI is near-optimal for binary evaluations
As the correlation $\rho$ is central to the sample efficiency factor $\tau$, we assume $b(m),p(m),q(m)\in(0,1)$ for the remainder of the text to exclude extreme cases in which $\rho(s(m),\tilde{s}(m))$ is undefined. It is now easy to upper bound the sample efficiency factor:
@@ 3.3 PPI is near-optimal for binary evaluations
The proof can be found in Appendix C.6.
This maximal variance reduction is only attained asymptotically: First, it requires unlimited proxy data, i.e. $N\to\infty$. Second, it assumes that the optimal $\lambda^{*}$ is known. In practice, $\lambda^{*}$ has to be estimated from data. This introduces additional variance and comes at the price of a bias of order $O(1/n)$ (Chaganty et al., 2018), if $\lambda^{*}$ is fit on the same data, or reduced sample size, if some data is set aside for estimating $\lambda^{*}$.
It is also worth noting that unless we have perfect correlation $\rho^{2}=1$, the sample efficiency factor $\tau(\hat{\theta}^{PP}_{\lambda^{*}})$ is finite. For any correlation bounded away from $1$, the proxy samples can only provide a constant factor improvement. Scaling them up without also scaling the number of ground truth samples has limited benefits. This raises the question whether we can find an unbiased estimator with lower variance. For black-box estimators that do not attempt to model the distribution of $(s(m),\tilde{s}(m))$ conditional on the prompt $x$, we show that the answer is no:
@@
Any unbiased estimator $\hat{\theta}$ for $\operatorname*{\mathbb{E}}s(m)$ that observes $n$ joint samples $(s(m),\tilde{s}(m))$ and $N$ independent proxy samples $\tilde{s}(m)$ fulfills the variance lower bound
@@
$\tau(\hat{\theta})\leq\tau(\hat{\theta}^{PP}_{\lambda^{*}})\leq\frac{1}{1-\rho(s(m),\tilde{s}(m))^{2}}\,.$
@@ 3.4 Limited gains for evaluating state-of-the-art models
With Theorem 5 at hand, we can analyze the effective sample size gains $\tau$ that unbiased estimators can achieve by making use of model judgments $\tilde{s}$. In this subsection, we focus on the evaluation of frontier models that outperform older models, including the models used to judge them. We use the evaluated model’s score $b(m)$ and the judge’s agreement rate $\operatorname*{\mathrm{AG}}(m)$ as indicators for the respective model’s capabilities. We focus on tasks for which performing and judging the task are of similar difficulty, such that these indicators are comparable. As a prime example, for binary classification, $\operatorname*{\mathrm{AG}}(m)$ equals the judge’s accuracy, while $b(m)$ equals the evaluated model’s accuracy.
@@
Assume $0.5\leq\operatorname*{\mathrm{AG}}(m)\leq b(m)$ and let $\Theta$ be the set of all unbiased estimators for $\operatorname*{\mathbb{E}}s(m)$ that observe $n$ joint samples $(s(m),\tilde{s}(m))$ and $N$ independent proxy samples $\tilde{s}(m)$. Then all estimators in $\Theta$ have a bounded sample efficiency factor:
$\tau_{\max}\coloneqq\max_{\hat{\theta}\in\Theta}\tau(\hat{\theta})\leq 2.$
@@ 3.5 High agreement rate is insufficient for sample efficiency gains
Theorem 6 holds for arbitrarily high levels of agreement. Therefore, the theorem shows that high agreement rates are not just insufficient for avoiding judge bias—they’re also insufficient for obtaining a meaningful sample efficiency factor $\tau$. In this subsection, we go a step further and show that the agreement $\operatorname*{\mathrm{AG}}(m)$ does not provide any lower bound on the sample efficiency factor $\tau$. Addressing this issue, we then propose a modified balanced agreement score and show that it can provide meaningful upper and lower bounds on $\tau$.
Our first proposition shows that agreement does not meaningfully constrain the squared correlation $\rho^{2}$, even when $\operatorname*{\mathrm{JB}}(m)$ is small. Correspondingly, in some cases access to proxy scores $\tilde{s}$ does not improve sample efficiency at all, even if $\tilde{s}$ turns out to be unbiased.
@@
$\operatorname*{\mathrm{JB}}(m)=\rho(s(m),\tilde{s}(m))^{2}=0.$
Proof.
We drop the dependence of our parameters on $m$ and set $q=1-b$ and $p=b$. Then $\operatorname*{\mathrm{JB}}=b(1-b)-(1-b)b=0.$ At the same time, as there is no judge bias, $\operatorname*{\mathbb{E}}\tilde{s}(m)=b=p$, such that $\rho(s(m),\tilde{s}(m))^{2}=0$. Lastly $\operatorname*{\mathrm{AG}}=b^{2}+(1-b)^{2}$, which can attain any value between $0.5$ and $1$. ∎
@@
In particular, the maximal sample efficiency factor $\tau_{\max}$ equals one in that case, such that despite access to $\tilde{s}$, no unbiased estimator has less variance than $\hat{\theta}^{GT}$. It might be surprising that there are cases in which we cannot make productive use of the proxy evaluations $\tilde{s}$ despite zero judge bias. However, without further assumptions, $\operatorname*{\mathrm{JB}}(m)$ can vary wildly between models, so that we cannot skip estimating it anew for each model. This is true even in cases where $\operatorname*{\mathrm{JB}}(m)$ turns out to be close to zero, post-hoc. Now, the estimator does not ”know” that $\tilde{s}$ has zero judge bias, but rather has to estimate $\operatorname*{\mathrm{JB}}(m)$ from joint observations $(s,\tilde{s})$. In the configuration from the proposition, estimating the judge bias $\operatorname*{\mathrm{JB}}(m)$ turns out to be as hard as directly estimating the real score $\operatorname*{\mathbb{E}}s(m)$.
3.6 Balanced agreement rate as an alternative
Intuitively, agreement fails to constrain the squared correlation $\rho^{2}$ because predictors without any signal, such as constant predictors, can achieve high agreement. This is analogous to how accuracy can be misleading for imbalanced classification problems. In classification, this is solved by focusing on balanced accuracy instead. By analogy, we introduce the balanced agreement of $\tilde{s}$ with $s$ as
$\operatorname*{\mathrm{BA}}(m)\coloneqq\frac{q(m)+p(m)}{2}.$

### v2

@@ 1.1 Our contributions
Figure 1: Model ranks on MMLU based on true labels compared to LLM labels in a semi-synthetic setting (a) and for the top-10 models on HELM by July 2024 (b). Using LLM labels heavily perturbs the ranking despite high judge accuracy.
@@
  • Our main result (Theorem 6) then shows that this correlation is small whenever the judge model performs worse than the evaluated model. In that case the best possible outcome is an effective doubling of the number of samples $s$, no matter how many samples of the proxy score $\tilde{s}$ are available (Corollary 7). This result remains true in a slightly weaker sense, when we allow for the proxy $\tilde{s}$ to be continuous rather than binary (Theorem 10).
@@
In addition, we show that the commonly reported agreement rate between the judge and the ground truth neither meaningfully restricts the judge bias, nor the sample size savings for debiasing methods.
@@ 2 Formal setup and motivation
In this section, we introduce the formal model we use to analyze the model-as-judge paradigm. We then provide empirical and theoretical evidence that judge bias can indeed cause model-as-judge to produce misleading results, even when judge models are quite accurate.
@@
2.1 Judge bias can strongly perturb model rankings
We argue that estimating the expected proxy score $\operatorname*{\mathbb{E}}\tilde{s}(m)$ instead of the real expected score $\operatorname*{\mathbb{E}}s(m)$ is problematic for benchmarking. In the spirit of Salaudeen & Hardt (2024), a benchmark has two purposes: quantifying model performance and ranking models. In turn, a useful LLM judge should accurately estimate each model’s performance, and induce the same ranking of evaluated models as the ground truth. For the first goal, the main quantity of interest is the judge bias, defined as
@@ 2.2 High agreement rate is insufficient for evaluation
The results from the previous subsection might appear surprising, considering the strong performance of the judge models: Lama-3 70B has $79\%$ accuracy on MMLU, while GPT-4 has $84\%$, and the models used in MT-Bench have $85\%$ agreement with expert annotators. Our next proposition explains how it is possible for model rankings to be strongly perturbed despite accurate judges. Before stating it, we need to formally define the agreement $\operatorname*{\mathrm{AG}}(m)$ between score $s$ and proxy score $\tilde{s}$ as
@@ 3 Debiasing LLM judgments
In the previous section, we demonstrated how biased LLM judges can lead to strongly perturbed model rankings. This motivates the need for debiasing methods. We begin this section by discussing a simple sufficient condition to guarantee asymptotically correct rankings: Ensure that our estimates of $\operatorname*{\mathbb{E}}s(m)$ are unbiased for all models $m$. Statistical bias correction methods can make use of a large number of proxy scores and a smaller number of gold standard labels to construct an unbiased estimator. We discuss one such method in detail, and show that it is essentially optimal in terms of estimator variance. Afterwards we prove our main result, i.e., that LLM judges offer limited benefits for evaluating state-of-the-art models.
@@ 3.1 Uniformly controlling judge bias ensures correct rankings
Proposition 3 implies that if we could guarantee sufficiently small estimator bias, a large amount of proxy samples $\tilde{s}$ would yield correct rankings with high probability. However, it is important for bias to be small for all evaluated models. In light of the worst-case results from Proposition 2, we cannot rely on this to be true when solely basing our estimates on the proxy scores $\tilde{s}$.
@@ 3.2 Background on Prediction Powered Inference
We next discuss how to leverage a small amount of ground truth samples $s$ to debias the proxy scores $\tilde{s}$. This yields an unbiased estimator $\hat{\theta}$ for $\operatorname*{\mathbb{E}}s$, and thus asymptotically correct model rankings.
@@
As a linear combination of unbiased estimators this remains unbiased, but we can now optimize over $\lambda$ to minimize estimator variance. We obtain the optimum value ${\lambda^{*}=\operatorname*{\mathrm{Cov}}(s(m),\tilde{s}(m))/\left(\left(1+\frac{n}{N}\right)\operatorname*{\mathrm{Var}}\tilde{s}\left(m\right)\right)}.$ At this $\lambda$, the estimator $\hat{\theta}^{PP}_{\lambda^{*}}$never increases variance compared to the classical ground truth estimator:
@@
In the following subsection, we discuss to what extent the PPI estimator $\hat{\theta}^{PP}_{\lambda^{*}}$ can reduce variance or equivalently improve sample efficiency, compared to the ground truth estimator $\hat{\theta}^{GT}$.
3.3 The sample efficiency factor $\tau$
As the variance of $\hat{\theta}^{GT}$ scales as $\Theta(1/n)$, using an estimator $\hat{\theta}$ with $r=\frac{\operatorname*{\mathrm{Var}}\hat{\theta}}{\operatorname*{\mathrm{Var}}\hat{\theta}^{GT}}$ is equivalent to increasing the ground-truth sample size by a factor of
$\tau(\hat{\theta})\coloneqq\frac{\operatorname*{\mathrm{Var}}\hat{\theta}^{GT}}{\operatorname*{\mathrm{Var}}\hat{\theta}}=\frac{1}{r}.$
We call $\tau(\hat{\theta})$ the sample efficiency factor of the estimator $\hat{\theta}$. It is our main quantity of interest. In order to ensure it is well-defined, we assume $b(m),p(m),q(m)\in(0,1)$ for the remainder of the text. Our next proposition provides an upper bound on the sample efficiency factor $\tau(\hat{\theta}^{PP}_{\lambda^{*}})$ of PPI, based on the squared Pearson correlation between $s$ and $\tilde{s}$:
@@ 3.3 The sample efficiency factor $\tau$
This upper bound is large whenever $\rho^{2}$ is large. However, the sample efficiency factor $\tau(\hat{\theta}^{PP}_{\lambda^{*}})$ is finite unless we have perfect correlation $\rho^{2}=1$. For any correlation bounded away from $1$, the proxy samples can only provide a constant factor improvement. Scaling them up without also scaling the number of ground truth samples has limited benefits. The proof can be found in Appendix C.7.
3.4 PPI is near-optimal for binary evaluations
The limited sample efficiency gains of PPI compared to the classical estimator raise the question whether we can find an unbiased estimator with lower variance. For black-box estimators that do not model the prompt-conditional distribution $\operatorname*{\mathbb{P}}[(s(m),\tilde{s}(m))|x]$, we show that the answer is no:
@@ 3.4 PPI is near-optimal for binary evaluations
Let $\Theta$ be the set of all unbiased estimators for $\operatorname*{\mathbb{E}}s(m)$ that observe $n$ joint samples $(s(m),\tilde{s}(m))$ and $N$ independent proxy samples $\tilde{s}(m)$. Then, any $\hat{\theta}\in\Theta$ fulfills the variance bound
@@
$\tau(\hat{\theta})\leq\max_{\hat{\theta}\in\Theta}\tau(\hat{\theta})=\tau(\hat{\theta}^{PP}_{\lambda^{*}})\leq\frac{1}{1-\rho(s(m),\tilde{s}(m))^{2}}\,.$
@@ 3.5 Limited gains for evaluating state-of-the-art models
With Theorem 5 at hand, we can analyze the best sample efficiency factor $\tau$ an unbiased estimator can achieve by making use of model judgments $\tilde{s}$. In this subsection, we focus on the evaluation of frontier models that outperform older models, including the models used to judge them. We use the evaluated model’s score $b(m)$ and the judge’s agreement rate $\operatorname*{\mathrm{AG}}(m)$ as indicators for the respective model’s capabilities. We focus on tasks for which performing and judging the task are of similar difficulty, such that these indicators are comparable. As a prime example, for binary classification, $\operatorname*{\mathrm{AG}}(m)$ equals the judge’s accuracy, while $b(m)$ equals the evaluated model’s accuracy.
@@
Assume $0.5\leq\operatorname*{\mathrm{AG}}(m)\leq b(m)$. Then,
$\tau_{\max}=\max_{\hat{\theta}\in\Theta}\tau(\hat{\theta})\leq 2.$
@@ 3.6 High agreement rate is insufficient for sample efficiency gains
Theorem 6 holds for arbitrarily high levels of agreement. Therefore, the theorem shows that high agreement rates are not just insufficient for avoiding judge bias—they’re also insufficient for obtaining a meaningful sample efficiency factor $\tau$. Our next proposition goes a step further and shows that the agreement $\operatorname*{\mathrm{AG}}(m)$ does not provide any lower bound on the sample efficiency factor $\tau$, even when the judge bias $\operatorname*{\mathrm{JB}}(m)$ is zero.
@@
$\operatorname*{\mathrm{JB}}(m)=\rho(s(m),\tilde{s}(m))^{2}=0\text{ and thus }\tau_{\max}=1.$
@@
We prove the proposition in Appendix C.8. It implies that in some cases despite access to $\tilde{s}$, no unbiased estimator has less variance than $\hat{\theta}^{GT}$. It might be surprising that there are cases in which we cannot make productive use of the proxy evaluations $\tilde{s}$ despite zero judge bias. However, without further assumptions, $\operatorname*{\mathrm{JB}}(m)$ can vary wildly between models, so that we cannot skip estimating it anew for each model. Now, even if $\tilde{s}$ has zero judge bias, the estimator does not ”know” that, but rather has to estimate $\operatorname*{\mathrm{JB}}(m)$ from joint observations $(s,\tilde{s})$. In the configuration from the proposition, estimating the judge bias $\operatorname*{\mathrm{JB}}(m)$ turns out as hard as directly estimating the real score $\operatorname*{\mathbb{E}}s(m)$.
3.7 Experiments
So far, we have shown that even optimal debiasing yields limited sample efficiency gains at the frontier (Section 3). In this section, we have a more detailed look into the sample efficiency gains on popular benchmarks that can be achieved by using current flagship models as judges.
Our experiments in this section focus on two settings: Multiple-choice question answering on MMLU (Hendrycks et al., 2021) and chatbot evaluation on MT-Bench (Zheng et al., 2024). Additional experiments on TruthfulQA can be found in Appendix B.2. In all cases, we calculate the empirical correlation $\rho(s,\tilde{s})$ between the score $s$ and the proxy $\tilde{s}$ to obtain and plot an upper bound for the sample efficiency factor $\tau_{\max}$ based on Theorem 5.
We first focus on MMLU, a multiple-choice question answering benchmark for the world knowledge of language models. We obtain model predictions from the HELM (Liang et al., 2023) leaderboard. We then aggregate all subtasks of MMLU into a single test set and set $s$ to equal the accuracy score $s(x,m)=\mathbb{I}(m(x)=y(x))$. Here $\mathbb{I}$ is the indicator function, $m(x)$ is the model’s answer and $y(x)$ is the correct answer. For model-based evaluations using $\tilde{m}$ as a judge, we replace the correct answer $y(x)$ by the judge’s prediction $\tilde{m}(x)$, setting $\tilde{s}(x,m)=\mathbb{I}(m(x)=m^{\prime}(x))$.
Figure 2: Best possible sample efficiency factor $\tau_{\max}$ according to Theorem 5, using different judges (colors) evaluating different models (x-ticks) on MMLU. Error bars show $90\%$ confidence intervals. Sample efficiency gains stay below two, unless SOTA models are used to evaluate weak models.
In the second setting, we evaluate model performance on MT-Bench, an arena-style benchmark designed to evaluate the ability of models to follow instructions. The benchmark comes with results for six models. For each pair of these models $m$ and $m^{\prime}$, both models are queried with the same prompt $x$. The test set for each model $m$ then consists of all triples of the form $(m,m^{\prime},x)$, for other models in the benchmark $m^{\prime}$ and prompts $x$. For each triple in the test set, an expert evaluator decides, whether model $m$ produced a better response than $m^{\prime}$. In that case, the score $s(x,m)$ for model $m$ is one, otherwise it is zero.  For the proxy score $\tilde{s}(x,m)$ we replace the expert’s answer by the GPT-4 response provided by Zheng et al. (2024). For more details, consider Appendix B.
Figure 3: Best possible sample efficiency factor $\tau_{\max}$ according to Theorem 5, using GPT-4-as-a-Judge, evaluating different models on MT-Bench. Error bars show $90\%$ confidence intervals. Sample efficiency get close to two in some cases, but consistently stay below that value.
Figures 2 and 3 show the optimal sample efficiency factor $\tau_{\max}$ on MMLU and MT-Bench respectively. The sample efficiency factor $\tau_{\max}$ consistently stays below the value of two suggested by Corollary 7, except when current flagship models are used to judge the significantly worse LLama2-7b on MMLU. This exception is not surprising, as in that case $\operatorname*{\mathrm{AG}}\gg b$ such that the assumptions for Corollary 7 are heavily violated. Interestingly, $\tau_{\max}$ stays below two in all other cases, even when stronger models like GPT-4 are used to evaluate weaker models like LLama3-70B. This suggests that the upper bound on $\tau_{\max}$ from Corollary 7 is fairly robust. We also note that $\tau_{\max}$ is often substantially smaller than two, especially on MMLU. This highlights that Corollary 7 provides a best-case upper bound rather than a guarantee that $\tau_{\max}$ will be close to two in practice.
4 Going beyond binary evaluations.
So far we have assumed the proxy score $\tilde{s}$ to be binary. In this section, we relax this assumption. By not forcing the judge to fully commit to a single answer, we hope to obtain a more useful proxy $\tilde{s}$. For example, in a Q&A task, we can make use of the uncertainty of a judge model $m^{\prime}$ by setting $\tilde{s}(x,m)=\operatorname*{\mathbb{P}}_{m^{\prime}(x)}(m(x))$ equal to the probability the judge $m^{\prime}$ assigned to the model answer $m(x)$.
In terms of analysis, the non-parametric joint likelihood of $(s,\tilde{s})$ resulting from non-binary proxies makes it exceedingly hard to prove an analogon to Theorem 5 for all unbiased estimators. Instead, we focus on the PPI estimator, $\hat{\theta}^{PP}_{\lambda^{*}}$. It is a natural choice of estimator, given its proven optimality in the special case of binary proxies $\tilde{s}$. In order to state a condition similar to $\operatorname*{\mathrm{AG}}\leq b$ from Theorem 6, we assume the proxy $\tilde{s}$ to be bounded in $[0,1]$ and define the soft agreement as
$\operatorname*{\mathrm{SO}}(\tilde{s})\coloneqq\operatorname*{\mathbb{E}}s\tilde{s}+(1-s)(1-\tilde{s}).$
For binary scores $\tilde{s}$, the soft agreement $\operatorname*{\mathrm{SO}}(\tilde{s})$ simply reduces to the agreement $\operatorname*{\mathrm{AG}}(\tilde{s})$. For non-binary proxy scores $\tilde{s}$, we interpret them as probability estimates, such that $\operatorname*{\mathrm{SO}}$ equals the estimated probability of the real score $s$. To make this interpretation valid, we focus on the recalibrated proxy ${R(\tilde{s})=\operatorname*{\mathbb{P}}(s=1|\tilde{s})}$. The following proposition suggests a way to generalize the condition $\operatorname*{\mathrm{AG}}\leq b$:
Proposition 9.
For any binary proxy $\tilde{s}$ such that $0.5\leq\operatorname*{\mathrm{AG}}\leq b$, we have $0.5\leq\operatorname*{\mathrm{SO}}(R(\tilde{s}))\leq b$.
Proposition 9 states that if the agreement $\operatorname*{\mathrm{AG}}(\tilde{s})$ of a binary proxy $\tilde{s}$ is below the evaluated model’s accuracy $b$, the same is true for the soft accuracy $\operatorname*{\mathrm{SO}}(R(\tilde{s}))$ of the recalibrated proxy $R(\tilde{s})$. Based on this, we use $\operatorname*{\mathrm{SO}}(R(\tilde{s}))\leq b$ to generalize the condition $\operatorname*{\mathrm{AG}}(\tilde{s})\leq b$ from Theorem 6, and prove: